Huang Automotive is presently operating at 75% of capacity. The company recently received an offer from a Korean truck manufacturer to purchase 22,500 units of a power steering system component for $197 per unit. Peter Wu, vice-president of sales, notes that although there will be an additional $3.00 shipping cost for each component, he thinks that accepting the order will get the company's "foot in the door" of an expanding international market.

To determine variable and fixed costs, Huang's accountant used the high-low method with the following production and cost information for the last two years:

202,000 units     227,000 units    
Direct material costs $16,564,000      $18,614,000     
Direct labor costs 5,555,000      6,242,500     
Overhead costs 23,014,000      24,439,000     
Selling and administrative costs 7,418,000      7,643,000     
Total costs $52,551,000      $56,938,500     
Total costs per unit $260.15      $250.83     



T.J. Chan, vice-president of engineering, feels that any new market should first show its profitability and that the $197 per unit offer is not only below the regular $260 selling price, but it's below the unit cost of the component. She also points out that there will be additional setup costs of $300,000 and that Huang will have to lease some special equipment for $215,000.

1. Using the high-low method to determine cost behavior, what would the expected profit be on the special order (use a negative sign for a loss)?
Answer #1
First step is to determine which cost is variable and which is fixed
Direct Material Cost
At 202,000 Units, Per unit Material Cost $                82.00 =16564000/202000
At 227,000 Units, Per unit Material Cost $                82.00 =18614000/227000
So Material Cost is same per year that means Material Cost is Variable Cost
Direct Labor Cost
At 202,000 Units, Per unit Labor Cost 27.5 =5555000/202000
At 227,000 Units, Per unit Labor Cost 27.5 =6242500/227000
So Labor Cost is same per year that means Labor Cost is also Variable Cost
Direct Overhead Cost
At 202,000 Units, Per unit Overhead Cost                  113.93 =23014000/202000
At 227,000 Units, Per unit Overhead Cost                  107.66 =24439000/227000
So Overhead Cost is not same per year that means Overhead Cost is Mixed Cost.
Variable Overhead Cost = Change in Cost / Change in Units
= ($24,439,000 - 23,014,000) / (227,000 – 202,000 Units)
Variable Overhead Cost = $57 Per Unit
Fixed Overhead Cost = Total Cost – Variable Cost
At 202,000 Units
= $23014000 – (202,000 Units * $57 per unit)
Fixed Overhead Cost = $11,500,000
Direct Selling and Adm. Cost
At 202,000 Units, Per unit Selling and Adm. Cost                    36.72 =7418000/202000
At 227,000 Units, Per unit Selling and Adm. Cost                    33.67 =7643000/227000
So Selling and Adm. Cost is not same per year that means Selling and Adm. Cost Cost is Mixed Cost.
Variable Selling and Adm. Cost = Change in Cost / Change in Units
= ($7643,000 - 7418,000) / (227,000 – 202,000 Units)
Variable Selling and Adm. Cost = $9.0 Per Unit
Fixed Selling and Adm. Cost = Total Cost – Variable Cost
At 202,000 Units
= $7418000– (202,000 Units * $9.0 per unit)
Fixed Selling and Adm. Cost = $5,600,000
If we will accept this order then we will only incur variable Cost, as Fixed Cost doesn’t change with change in Production or Sales.
Cost
Sales (22,500 Units * 197) 4432500
Less: Cost
Existing Cost
Material (22,500 Units * 82) 1845000
Labor (22,500 Units * 27.5) 618750
Variable Overhead (22,500 Units * 57) 1282500
Variable Selling and Adm. (22,500 Units * 9.0) 202500
New Cost
Shipping Cost (22,500 Units * 3.0) 67500
Setup Cost 300000
Lease Cost 215000
Total Cost 4531250
Net Income / (Loss) -98750

There would be loss of $98,750

Note:
There are some new cost which we will consider in Decision Making
Conclusion
We should not accept this order as it will decrease our Net Income by $98,750

Please give positive rating.